OpenStudy (anonymous):

Find the average rate of change of f from 0 to 7pi/12. f(x)=tan(2x)

OpenStudy (anonymous):

(4sqrt3)/(7pi)

OpenStudy (anonymous):

Correct?

myininaya (myininaya):

\[\frac{\tan(2 \cdot \frac{ 7 \pi}{12})-\tan(2 \cdot 0)}{\frac{7 \pi }{12}-0}\]

myininaya (myininaya):

let me find out by simplifying this

OpenStudy (anonymous):

yes it is correct, since it starts at (0,0)

myininaya (myininaya):

\[\frac{\tan(\frac{7 \pi}{6})}{\frac{7 \pi}{12}}\]

myininaya (myininaya):

\[\frac{12}{7 \pi} \cdot \tan(\frac{ 7 \pi}{6} )\]

OpenStudy (anonymous):

oops i was wrong \[\tan(\frac{7\pi}{6})=\frac{1}{\sqrt{3}}\]

myininaya (myininaya):

\[\frac{12}{7 \pi} \cdot \frac{\sin(\frac{7 \pi}{6})}{\cos(\frac{7 \pi}{6})}\] \[\frac{12}{7 \pi} \cdot \frac{\sin( \pi +\frac{\pi}{6})}{\cos( \pi +\frac{\pi}{6})}\] \[\frac{12}{ 7 \pi} \cdot \frac{\frac{-1}{2}}{\frac{-\sqrt{3}}{2}}=\frac{12}{7 \pi} \cdot \frac{1}{\sqrt{3}}\] lol sorry i was having a find an error in my latex

OpenStudy (anonymous):

No worries